OpenStudy (kirbykirby):

Inverse Laplace transform: \[\mathcal{L}^{-1} \left( \frac{5}{s(s^2+4s+5)} \right) \]

OpenStudy (accessdenied):

Since we have that grouping of two factors in the bottom, I think a good first step is to break it up by partial fractions and then use the fact that: \(\mathcal{L}^{-1} \left[F + G\right] = \mathcal{L}^{-1} \left[F\right] + \mathcal{L}^{-1} \left]G \right]\)

OpenStudy (kirbykirby):

\[ 5*\mathcal{L}^{-1} \left( \frac{1/5}{s}-\frac{s}{s^2+4s+5}-\frac{4}{s^2+4s+5} \right) \]

OpenStudy (kirbykirby):

oops \[ 5*\mathcal{L}^{-1} \left( \frac{1/5}{s}-\frac{-s/5}{s^2+4s+5}-\frac{4/5}{s^2+4s+5} \right) \]

OpenStudy (kirbykirby):

\[ \mathcal{L}^{-1} \left( \frac{1}{s} \right) - \mathcal{L}^{-1} \left( \frac{s}{(s+2)^2+1} \right)- 4*\mathcal{L}^{-1} \left( \frac{1}{(s+2)^2+1} \right) \]

OpenStudy (kirbykirby):

\[ 1 - e^{-2t}cos(t)- 4e^{-2t}sin(t) \] ??

OpenStudy (accessdenied):

Hmm... I think the second term is not correct \( \displaystyle \mathcal{L}^{-1} \left[ \frac{s-a}{(s-a)^2+ b^2} \right] = e^{at} \cos bt \) We need b=1 and a=-2, but the numerator is only s and not s+2. Instead, we could pull out the extra +2 we need from that last term. The last term won't be affected besides a smaller coefficient. (2 instead of 4).

OpenStudy (accessdenied):

\( \displaystyle -\mathcal{L}^{-1} \left[\frac{s+4}{(s + 2)^2 + 1} \right] = -\mathcal{L}^{-1} \left[ \frac{s+2}{(s+2)^2 + 1} + \frac{2}{(s + 2)^2 + 1} \right] \)

OpenStudy (kirbykirby):

oh that's interesting. I thought I could use the simpler cosine transform , \[\mathcal{L}[cos(bt)]=\frac{s}{s^2+b^2}\] and use the shift theorem.

OpenStudy (kirbykirby):

oh because there's an "s" on top as well.. -_-

OpenStudy (accessdenied):

You could, although we have to be careful b/c the shift changes all s \(\to\) s-a, so that'd be both the s in the numerator and denominaotr becoming \(s-a\) But that will also work. :)

OpenStudy (kirbykirby):

I get it now :)

OpenStudy (kirbykirby):

lol sorry for making you write all that I just noticed it now

OpenStudy (accessdenied):

Oh, sorry, I think my internet just played a trick on me. lol I didnt see that last msg. :) It's fine, and glad you got it now! :D

OpenStudy (kirbykirby):

thank you soo much :)!!

OpenStudy (accessdenied):

You're welcome!